OpenStudy (anonymous):

ASAP PLEASE!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! Explain, in complete sentences, the process of using Cramer’s Rule to solve the following system of equations. Provide the solution to the system. 2x + 7y = 2 3x + 8y = 2

OpenStudy (anonymous):

i can solve it but not using cramers rule solve for one variable in one and plug it into thhe other equation. x = -2/5, y = 2/5

OpenStudy (anonymous):

i have to have it using the rule

OpenStudy (anonymous):

http://www.purplemath.com/modules/cramers.htm

OpenStudy (anonymous):

does that help?

OpenStudy (anonymous):